OpenStudy (anonymous):

2 over 6 - radical 3

OpenStudy (lgbasallote):

\[\frac{2}{6-\sqrt 3}\] multiply by conjugate \[\frac{2}{6-\sqrt 3} \times \frac{6+\sqrt 3}{6+\sqrt 3} \implies \frac{2(6+\sqrt 3)}{(6 + \sqrt 3)(6 - \sqrt 3)} \implies \frac{12 + 2\sqrt 3}{6^2 - (\sqrt 3)^2} \implies \frac{12 + 2\sqrt 3}{36 - 3}\] do you get it? or do you need more help?

OpenStudy (anonymous):

I Understand Thank You Soo Much ^.^

OpenStudy (lgbasallote):

myy pleasure :D